OpenStudy (anonymous):

Solve the logarithmic equation. log 5 + log x – log 7 = 2

OpenStudy (anonymous):

by log do you mean ln()?

OpenStudy (anonymous):

log (5xX / 7) =2 5X/7 = 10^2 5X = 7x10^2 X = ? try it :3

OpenStudy (anonymous):

lemme know if u didn't get it :D i'll explain it ;)

OpenStudy (whpalmer4):

Remember that adding the logs of two numbers is the same as taking the log of their product: \[\log a + \log b = \log (a*b)\] and subtracting two logs is the same as taking the log of their quotient:\[\log a - \log b = \log (a/b)\]That is the key to understanding @feezul_93's first step.

OpenStudy (anonymous):

yeap, exactly, but i planned to let his brain work 1st, I bet he learnt it before :p

OpenStudy (anonymous):

finally my question loaded lemme work that

OpenStudy (anonymous):

@androidonyourface we answered your question d :)

OpenStudy (whpalmer4):

I would have led with the principles, and let him try to apply them, rather than just figuring out what you did...

OpenStudy (whpalmer4):

I agree that he has probably seen the principles before, but easy to forget!

OpenStudy (anonymous):

i used to learnt the concept by workflow, and remembered these thingy when people making fun of me, lol :p

OpenStudy (anonymous):

I got it thanks sorry for some reason the question goes blank then I gotta refresh to see yalls comments

OpenStudy (anonymous):

but no doubt, u're doing fine though ;)

OpenStudy (whpalmer4):

What's your final answer?

OpenStudy (anonymous):

140 :)